OpenStudy (anonymous):

Solve for x. 2 • 3 – 2 • 11 = 2(x – 11) A. 9 B. 3 C. 2 D. 6

OpenStudy (bibby):

PEDMAS is your order of operations. Start by distributing the 2 on the right hand side

OpenStudy (haseeb96):

2.3-2.11=2(x-11) 6-22=2x-22 -16+22=2x 6=2x x=3 it is the absolutly correct answer

OpenStudy (haseeb96):

I means C is the correct answer